### v3

# High-throughput polarization-independent spatial mode shaping with mode-selective photonic lanterns

Abstract
Coherent interference in multimode photonic systems offers a transformative approach for scalable, high-fidelity optical control, enabling precise beam shaping, efficient power delivery, and advanced signal processing through single-mode performance. Current techniques are often constrained to a single polarization state of light, system complexity, and the fabrication challenges associated with precise waveguide structures. Here, we present a high-efficiency coherent combination scheme employing a commercial mode-selective photonic lantern supporting three spatial channels, achieving Gaussian mode interference at multiple locations along the multimode core with up to 100% mode conversion efficiencies. By integrating a Faraday reflecting fibre component and piezo-phase shifters with the photonic lantern channels, we demonstrate stable and adaptive mode manipulation without the need for complex adaptive optics. These findings highlight the potential of photonic lanterns as robust platforms for controlled mode interference, paving the way for their deployment in next-generation optical communication, beam shaping, and beam-forming systems.
preprint: AIP/123-QED
Optical fibres are fundamental components in modern photonic systems, enabling the transmission of optical signals across a wide range of applications, including long-haul optical communicationsdoi:10.1049 ; Fatome:10 , quantum key distributionarticleqkd1 ; 20.044006 , and imagingscience.abl3771 ; Li:00 . While single-mode fibres (SMFs) enable higher bandwidth for signal transmissionarticle222 , their small core size reduces coupling efficiency, requiring precise alignment and introducing instrumental complexity1128517 . One approach to improve coupling efficiency is to use larger-core multimode fibres (MMFs)601675 ; 9277524 . However, a fundamental limitation of MMFs in signal transport applications is modal dispersion, arising from the complex propagation dynamics of light within the larger coreKovaevi2017ARA ; 2010xzc . Despite the challenges of signal coupling, modern optical communication systems continue to favor SMFs for their high bandwidth capacity and minimal signal dispersion. An ideal optical fibre system would combine the high multiplexed coupling efficiency of MMFs with the bandwidth and signal-processing advantages of SMFs. Significant research efforts have focused on mitigating modal dispersion and signal broadening in MMFs through advanced wavefront shapingarticlebnm ; articlevgf ; 1.5136334 ; 9420070 and adaptive optics-based beam combination techniquesMORI2013132 ; articlecdt ; Castillo:25 . However, despite these advances, achieving multimode light coupling and single-mode conversion in waveguide systems still relies on complex adaptive optical componentsChen:15 ; refId0 .
With the increasing demand for space-based quantum communication, a hybrid multimode-to-single-mode (MMF-SMF) system offers compelling advantages for quantum key distribution (QKD) receivers by enabling efficient multiplexed light collection while preserving single-mode performance, critical for minimizing modal dispersion and enhancing timing precision in photon detection. In QKD, a promising solution for multimode light collection and efficient reformatting of multimode to single-mode signals is the use of photonic lanterns (PLs)Leon-Saval:10 ; Birks:15 . With a multimode input and multiple single-mode (or few-mode) outputs, PLs have proven to be excellent candidates for multiplexed light collection, detection, and processing. In addition to specific applications in time-stretch imagingarticleHKC , compressive imagingarticleCD , and optical coherence tomographyarticleOCT , the PLs now offer significant advancement in QKD receiver performance and represent a promising alternative to complex adaptive optics systems in free-space optical communications and QKDBillault:25 ; beraza2025 ; Chandrasekharan:24 .
Figure 1: Experimental system used for accousto-optic coherent mode combination system. The red arrow denotes the forward-propagating optical signal, whereas the black arrow indicates the trajectory of the back-reflected light. Inset - PL mode profiles obtained at the MM input when coupling light from the PL output fibres, (i)LP01 mode (ii) LP 11a mode, (iii) LP11b mode. P – polarizer, PBS – polarizing beamsplitter, BS- beamsplitter, QWP – quarter waveplate, PLO- photonic lantern output, FDL- fibre delay line, PE, piezoelectric, FRM- Faraday rotating mirror. Scale bar, 0.5 mm.
In this paper, we report a proof-of-concept experimental demonstration of piezoelectric (PE) phase-shifter coherent mode recombination, achieving near-unity (100%) mode conversion efficiency within a commercially available three-mode PL. Our approach leverages acousto-optic modulation combined with a Faraday reflecting fibre to dynamically and precisely control the relative phases of the individual guided modes. This phase tuning enables deterministic spatial manipulation of the resulting interference pattern, allowing the targeted single-mode Gaussian-like output profile to be positioned at arbitrary locations across the multimode core of the PL. This capability represents a significant advancement in integrated photonic systems, offering scalable and polarization-independent control of mode interference without reliance on complex adaptive optics. Our results underscore the potential of PE phase-shifter coherent control schemes to enable robust, high-fidelity mode conversion for advanced beam shaping, signal processing, and next-generation photonic communication technologies.
[…]
To demonstrate the state-of-the-art, fibre delay lines (FDLs) and fibre Faraday reflecting mirrors (FRMs) were incorporated at each PL output with equal lengths, enabling both reflection from the FRMs and precise phase control of the back-reflected light. The FDLs were configured to allow path length adjustment by varying the relative positions of the input and output fibre facets using a linear translation stage with a 60 mm travel range. In our implementation, fine phase modulation was achieved by introducing PE actuators within the delay lines, where small path length variations were controlled by applying voltage signals to the PE stacks. Three PE stacks were used to independently adjust the optical path length—and thus the phase—at each of the three PL outputs.
Figure 2: (A). 20 random mode interference profiles obtained at the MM end of the PL for various piezo voltage combinations. The frames are extracted from the on-the-fly video file during the voltage scan. As seen, the mode energy is concentrated in Gaussian-like profiles on various voltage combinations. The dotted circle outlines the approximate core region of the multimode end. All images cover the same spatial area, scale bar, 0.5 mm. (B) Respective mode energies in the image frames, quantified by summing the pixels within the core region. (C) 3D visualization of 6 random images with smoothed surface rendering and corresponding images projected on YZ wall, capturing the diverse spatial evolution of the light field upon accousto-optic scan. The final frame exhibits a Gaussian-like intensity distribution, indicative of effective multimode beam shaping and high-fidelity conversion to a desired single-mode spatial profile.
Figure 3: (A) Representative mode intensity profiles extracted from the recorded video sequence, illustrating key interference outcomes. Subfigure (a) presents the image frame with the minimum photon flux, subfigure (b) highlights a highly non-Gaussian interference state, subfigures (c-f) present image frames with distinct instances of Gaussian-like mode interference patterns. Subfigure (f)- the mode profile associated with the maximum photon flux. Scale bar, 0.5 mm. (B) Total photon counts for each of the six frames, calculated by integrating pixel intensities within the defined MM core region. Corresponding mode profiles are shown as insets for direct comparison between spatial distribution and total energy. The bar plot shows the energy conversion efficiencies of the six mode profiles (subfigures (a–f)), calculated by normalizing their photon fluxes to that of subfigure (f), which exhibits the highest photon flux in the entire video. This comparison illustrates the variation in energy conversion efficiency across different spatial interference patterns.
Light was injected into the MM end of the PL, and the relative output intensities were measured as 46%, 29%, and 25% for PLO1, PLO2, and PLO3, respectively. This relative energy distribution was quantified by selectively measuring the signal intensities detected at each of the PL outputs. Once the coupling conditions were carefully established, the effect of phase-induced light coupling of the returned signals at the PL outputs was investigated. A MATLAB script was used to interface the voltage generator to send and modulate the voltage for PE stacks for controlled phase changes of the back-reflected light from the PL outputs.
The PE stacks, designed to achieve a maximum displacement of 150 $\mu$m, were used to modulate the phase of the incoming light as a function of the applied voltage. A series of voltage values was applied to the PE stacks in a programmable sequence, inducing controlled phase shifts across the PL output delay lines. These phase shifts produced varying interference conditions, which were highly sensitive to even small changes in phase. The phase-induced interference effects were monitored on-the-fly using the camera positioned at the reflection arm of the PBS. To fully capture the mode profile changes at the MM end, the voltage sequence was set to update every 1 second, while the camera was operated at a frame rate of 5–10 frames per second. The entire process—including on-the-fly voltage adjustments and image capture—was recorded using the FastStone Capture screen recorder software. This ensured that the voltage settings and corresponding interference patterns could be replicated, which is essential for future mode combination strategies.
Figure 2A displays 20 random representative image frames (background noise corrected) extracted from a continuous voltage scan, selected to showcase the diversity of optical field distributions resulting from dynamic phase modulation at the PL output. These frames reveal a broad range of interference conditions between the spatial channels, including multiple instances of constructive and destructive interference. Notably, certain frames exhibit strong spatial localization with Gaussian-like intensity profiles, indicating coherent mode superposition with high modal overlap. Such energy localization enables precise and efficient multimode-to-single-mode conversion, which is particularly beneficial for single-photon QKD receivers, where spot size directly impacts the achievable key rateqtc2.12091 .
To provide a quantitative assessment of the interference dynamics, the total photon count in each frame was computed by summing the pixel intensities within the defined multimode core region. The resulting values, plotted in Figure 2B, exhibit significant fluctuations corresponding to the evolving interference states. Peaks in the total count trace denote constructive interference events with efficient modal combination, while dips indicate destructive interference and phase mismatch between channels. This analysis underscores the system’s capacity for real-time phase-sensitive control over mode interference patterns at the output facet.
Figure 2C presents a sequence of smoothed 3D surface plots of 6 random output intensity profiles from the PL, captured while scanning the voltage applied to the acousto-optic modulator. Each plot is normalised and accompanied by the corresponding 2D intensity projection on the yz-plane, illustrating the evolving interference patterns at the multimode output as the phase of the back-reflected signal is tuned via the fibre-integrated Faraday reflector. The surface profiles highlight the spatial evolution of mode combinations across the lantern output, revealing how varying voltage values modulate the resultant intensity structure. This figure visually demonstrates the platform’s ability to achieve controlled and reconfigurable modal outputs through all-fibre, polarization-independent tuning.
A key novelty of our experimental system is its ability to dynamically position Gaussian-like mode profiles at different locations across the MM end. Notably, for this set of measurements, two distinct locations of Gaussian-like interference were observed as shown in Figure 3A(e, f), achieving remarkable energy conversion efficiencies of 95% and 100%, respectively. These high-efficiency spots demonstrate the effectiveness of coherent mode control within the PL architecture. The ability to relocate and stabilise such modal interference patterns without active wavefront shaping underscores the scalability and practical robustness of our polarization-independent system for advanced beam-shaping and signal-processing applications.
To conclude, we have demonstrated an all-fibre, polarization-independent platform capable of real-time, reconfigurable modal control at the input MM end of a PL, enabling multimode light injection and single-mode light enhancement. By scanning the voltage applied to an integrated acousto-optic modulator, we achieved dynamic phase modulation of back-reflected signals, resulting in a wide range of interference conditions between spatial channels. Our analysis revealed both constructive and destructive modal interference, with certain configurations yielding strongly localised, Gaussian-like intensity profiles. Crucially, these profiles were not only repeatable but also tunable to distinct spatial positions across the multimode output, achieving energy conversion efficiencies of up to 100%. This highlights the system’s ability to precisely manipulate modal content through coherent interference, without the need for active wavefront shaping or polarization control. The results underscore the robustness, scalability, and practical relevance of this approach for advanced photonic applications, including spatial mode switching, adaptive optics, and coherent beam combination.
ACKNOWLEDGMENTS
The authors thank Innovate UK (10005967); Royal Academy of Engineering RF\201718\1746); Engineering and Physical Sciences Research Council (EP/T001011/1) for funding support.
AUTHOR DECLARATIONS
Conflict of Interest
The authors declare no conflict of interests.
DATA AVAILABILITY
The data that support the findings of this study are openly available in the Heriot-Watt University PURE research data management system.